# Complexity reduction of physical models: an equation-free approach by means of scaling

### v1

@@ Appendix B Theorems
By employing the notation (1) in (A.1) (Property A.1), one has
$\lambda_{i}=\tilde{\kappa}_{i}\,\tilde{\theta}_{1}^{c_{i,1}}\,\tilde{\theta}_{2}^{c_{i,2}}\,\cdots\,\tilde{\theta}_{N_{\mathsf{x}}}^{c_{i,N_{\mathsf{x}}}}=\kappa_{i}[\tilde{\kappa}_{i}]\,\theta_{1}^{c_{i,1}}[\tilde{\theta}_{1}]^{c_{i,1}}\,\theta_{2}^{c_{i,2}}[\tilde{\theta}_{2}]^{c_{i,2}}\,\cdots\,\theta_{N_{\mathsf{x}}}^{c_{i,{N_{\mathsf{x}}}}}[\tilde{\theta}_{N_{\mathsf{x}}}]^{c_{i,{N_{\mathsf{x}}}}},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.3)
$[\lambda_{i}]=[\tilde{\kappa}_{i}]\,[\tilde{\theta}_{1}]^{c_{i,1}}\,[\tilde{\theta}_{2}]^{c_{i,2}}\,\cdots\,[\tilde{\theta}_{N_{\mathsf{x}}}]^{c_{i,{N_{\mathsf{x}}}}}=1,\quad\forall i=1,\dots,N_{\mathsf{l}},$ (B.4)
it follows from (A.1) and (B.3) that
$\lambda_{i}=\kappa_{i}\,\theta_{1}^{c_{i,1}}\,\theta_{2}^{c_{i,2}}\,\cdots\,\theta_{N_{\mathsf{x}}}^{c_{i,{N_{\mathsf{x}}}}}\in(0,\infty),\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.5)
Then, (B.5) allows writing the function $C_{U}$ as
$C_{U}(\vec{\rho})=\sum_{i\in U}\left(\log_{10}(\kappa_{i})+\sum_{k=1}^{N_{\mathsf{x}}}\,c_{i,k}\,\rho_{k}\right)^{2},\quad\forall\vec{\rho}\coloneqq(\rho_{1},\dots,\rho_{N_{\mathsf{x}}})\in\mathbb{R}^{N_{\mathsf{x}}},$ (B.6)
where $\rho_{k}\coloneqq\log_{10}(\theta_{k})$ for any $k=1,\dots,N_{\mathsf{x}}$. The twice-differentiable function $C_{U}(\vec{\rho})$ (B.6) is convex because its Hessian matrix is positive semi-definite $\forall\vec{\rho}\in\mathbb{R}^{N_{\mathsf{x}}}$. Then, as guaranteed by Theorem 2.5 in [19], the global minimizers of $C_{U}$ can be found by imposing $\partial_{\rho_{j}}C_{U}=0$, $\forall j=1,\dots,N_{\mathsf{x}}$, yielding to the linear system (B.2) of $N_{\mathsf{x}}$ equations with $N_{\mathsf{x}}$ unknowns $\rho_{1},\dots,\rho_{N_{\mathsf{x}}}\in\mathbb{R}$.
Theorem B.2 algebraically formalizes a criterion for dimensionlessness of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$.
Theorem B.2.
Given Property A.6 of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, any quantity of the form
$\tilde{p}_{1}^{z_{1}}\,\tilde{p}_{2}^{z_{2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{z_{N_{\mathsf{p}}}},\quad\mbox{with}\quad z_{1},z_{2},\dots,z_{N_{\mathsf{p}}}\in\mathbb{R},$ (B.7)
is dimensionless if and only if the vector $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}$ of exponents in (B.7) belongs to the null space (kernel) of the matrix
$\textbf{M}\coloneqq\begin{pmatrix}m_{1,1}&m_{1,2}&\cdots&m_{1,N_{\mathsf{p}}}\\ m_{2,1}&m_{2,2}&\cdots&m_{2,N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ m_{N_{\mathsf{u}},1}&m_{N_{\mathsf{u}},2}&\cdots&m_{N_{\mathsf{u}},N_{\mathsf{p}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{u}}\times N_{\mathsf{p}}},$ (B.8)
where the entries of M (B.8) are the exponents in (A.4) specified by Property A.6.
@@
Any quantity of the form (B.7) is unitless if and only if its dimensions are
$[\tilde{p}_{1}]^{z_{1}}\,[\tilde{p}_{2}]^{z_{2}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{z_{N_{\mathsf{p}}}}=1,$ (B.9)
or, equivalently, by using (A.4) (Property A.6),
$[\tilde{p}_{1}]^{z_{1}}\,[\tilde{p}_{2}]^{z_{2}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{z_{N_{\mathsf{p}}}}=\mathrm{u}_{1}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{1,j}\,z_{j}}\,\mathrm{u}_{2}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{2,j}\,z_{j}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{N_{\mathsf{u}},j}\,z_{j}}=1.$ (B.10)
From independence of $\mathrm{u}_{1},\dots,\mathrm{u}_{N_{\mathsf{u}}}$ assured by Property A.6, (B.10) is equivalent to
$\sum_{j=1}^{N_{\mathsf{p}}}m_{i,j}\,z_{j}=0,\quad\forall i=1,\dots,N_{\mathsf{u}},\quad\mbox{i.e.,}\quad\textbf{M}\,\vec{z}=\vec{0},$ (B.11)
for M defined in (B.8) and $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}$. In other words, any quantity of the form (B.7) is dimensionless if and only if the vector $\vec{z}$ of exponents in (B.7) belongs to the null space (kernel) of matrix M (B.8), i.e., $\vec{z}\in\mathrm{ker}(\textbf{M})$.
Theorem B.3 determines the minimal number of independent dimensionless parameters and indicates a way to compute them in terms of the provided physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$.
Theorem B.3.
Given Properties A.4 to A.6 of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, there exist only $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ of the form (B.7), with M as in (B.8). Such parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ are identified by the vectors composing a basis of the null space (kernel) of the matrix M (B.8):
$\mbox{with}\quad\vec{b}_{1},\vec{b}_{2},\dots,\vec{b}_{N_{\mathsf{0}}}\quad\mbox{constituing a}\quad\mathrm{basis}(\mathrm{ker}(\textbf{M})),$ (B.12)
Proof.
First, we show that Properties A.4, A.5 and A.6 imply that the parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) are dimensionless, strictly positive and independent. Then, we conclude the proof of Theorem B.3 by demonstrating that any dimensionless quantity of the form (B.7) can be written in terms of such $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, as this means that there are no further independent quantities one can consider.
Given the matrix M (B.8) of exponents in (A.4) (Property A.6), any basis of $\mathrm{ker}(\textbf{M})$ is composed by $N_{\mathsf{0}}\coloneqq\mathrm{dim}(\mathrm{ker}(\textbf{M}))=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ independent vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}\in\mathrm{ker}(\textbf{M})\subseteq\mathbb{R}^{N_{\mathsf{p}}}$. The corresponding $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) are dimensionless since $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}\in\mathrm{ker}(\textbf{M})$, as guaranteed by Theorem B.2 assuming Property A.6.
Then, by employing the notation (1) in (B.12), one has
@@
$[\pi_{k}]=[\tilde{p}_{1}]^{b_{k,1}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{b_{k,{N_{\mathsf{p}}}}}=1,\quad\forall k=1,\dots,N_{\mathsf{0}},$ (B.14)
and $p_{1},\dots,p_{N_{\mathsf{p}}}\in(0,\infty)$ (Property A.5), it follows the strict positiveness of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12), i.e.,
@@
Moreover, such $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) are independent since
@@
On the one hand, $a_{1}=a_{2}=\cdots=a_{N_{\mathsf{0}}}=0$ trivially implies that $\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}}=1$. On the other hand, given $\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}}=1$, it follows from (B.12) that
@@
which implies, from independence of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ assured by Property A.4,
@@
As $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ constitute a basis of $\mathrm{ker}(\textbf{M})$, they are linearly independent vectors. Thus, (B.18) gives $a_{1}=a_{2}=\cdots=a_{N_{\mathsf{0}}}=0$ and (B.16) holds, showing that $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) are independent.
Finally, any dimensionless quantity of the form $\tilde{p}_{1}^{z_{1}}\,\tilde{p}_{2}^{z_{2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{z_{N_{\mathsf{p}}}}$, with $z_{1},z_{2},\dots,z_{N_{\mathsf{p}}}\in\mathbb{R}$, is such that $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathrm{ker}(\textbf{M})$, as guaranteed by Theorem B.2 assuming Property A.6. Then, the vector $\vec{z}$ can be written as a linear combination of the vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ composing a basis of $\mathrm{ker}(\textbf{M})$. By using (B.12), it follows that
@@
for some $a_{1},\dots,a_{N_{\mathsf{0}}}\in\mathbb{R}$. In conclusion, any dimensionless quantity of the form (B.7) can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) only. ∎
Theorem B.4 states that the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written in terms of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) only.
Theorem B.4.
Given Properties A.1 to A.6 and $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12), each of the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, i.e.,
$\lambda_{i}=\pi_{1}^{\alpha_{i,1}}\,\pi_{2}^{\alpha_{i,2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},\quad\alpha_{i,k}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall k=1,\dots,N_{\mathsf{0}},$ (B.20)
where the exponents $\alpha_{i,k}$ in (B.20) are unknowns of scaling procedures.
@@
As it does for $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) (see the proof of Theorem B.3), Property A.5 ensures that the numerical values $\kappa_{i}$ and $\theta_{j}$ of all $\tilde{\kappa}_{i}$ in (A.2) and $\tilde{\theta}_{j}$ in (A.3) are strictly positive. Then, it is possible to plug (A.2) and (A.3) into (A.1) (see Properties A.1, A.2 and A.3). This uncovers that each of the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only, i.e., in the form (B.7):
$\exists\,\vec{l}_{i}\coloneqq(l_{i,1},l_{i,2},\dots,l_{i,N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}\quad\mbox{such that}\quad\lambda_{i}=\tilde{p}_{1}^{l_{i,1}}\,\tilde{p}_{2}^{l_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{l_{i,N_{\mathsf{p}}}},$ (B.21)
@@
Each of the coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (B.21) is dimensionless and, thus, Theorem B.2 guarantees, under Property A.6, that each of the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ of exponents in (B.21) belongs to the kernel of the matrix M (B.8), i.e., $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}\in\mathrm{ker}(\textbf{M})$. If one considers the basis’ vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12) (Theorem B.3) assuming Properties A.4, A.5 and A.6, any vector $\vec{l}_{i}$ (B.21) belonging to $\mathrm{ker}(\textbf{M})$ can be written (in a unique way) as a linear combination of such $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12). Then, it follows
@@
By plugging (B.22) into (B.21) and, then, using (B.12) (Theorem B.3) assuming Properties A.4, A.5 and A.6, it follows that each of the coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) only, i.e.,
$\lambda_{i}=\tilde{p}_{1}^{l_{i,1}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{l_{i,N_{\mathsf{p}}}}=\tilde{p}_{1}^{\sum_{k=1}^{N_{\mathsf{0}}}\alpha_{i,k}\,b_{k,1}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\sum_{k=1}^{N_{\mathsf{0}}}\alpha_{i,k}\,b_{k,N_{\mathsf{p}}}}=\pi_{1}^{\alpha_{i,1}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},$ (B.23)
being $l_{i,j}$ and $b_{k,j}$ the $j$-th components of the vectors $\vec{l}_{i}$ and $\vec{b}_{k}$, respectively, for any $i=1,\dots,N_{\mathsf{l}}$, $j=1,\dots,N_{\mathsf{p}}$ and $k=1,\dots,N_{\mathsf{0}}$. We remark that the values of the exponents $\alpha_{i,k}$ in (B.23) are not explicitly known and, in this sense, we state that they are unknowns of scaling procedures. ∎
Theorem B.5 offers a way to obtain the unknown exponents $t_{i,j}$ in (A.3).
Theorem B.5.
Given Properties A.1 to A.6, the vector
$\vec{t}\coloneqq\begin{pmatrix}\vec{t}_{1}\\ \vec{t}_{2}\\ \vdots\\ \vec{t}_{N_{\mathsf{x}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\,N_{\mathsf{x}}},\quad\mbox{with}\quad\vec{t}_{j}\coloneqq\begin{pmatrix}t_{1,j}\\ t_{2,j}\\ \vdots\\ t_{N_{\mathsf{p}},j}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}},$ (B.24)
of unknown exponents in (A.3) must satisfy the linear system
$\textbf{Y}\,\vec{t}=\vec{\gamma},\quad\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\gamma}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}},$ (B.25)
where the entries of Y are explicitly known real numbers, while the entries of $\vec{\gamma}$ are explicit affine functions of the exponents $\alpha_{i,k}$ in (B.20), as specified by (B.38).

### v2

@@ Appendix A Proof of Theorem 1
Any quantity of the form (19) is unitless if and only if its dimensions are
$\tilde{p}_{1}]^{z_{1}}\,[\tilde{p}_{2}]^{z_{2}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{z_{N_{\mathsf{p}}}}=1,$ (A.1)
or, equivalently, by using (17) (Assumption˜6),
$\tilde{p}_{1}]^{z_{1}}\,[\tilde{p}_{2}]^{z_{2}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{z_{N_{\mathsf{p}}}}=\mathrm{u}_{1}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{1,j}\,z_{j}}\,\mathrm{u}_{2}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{2,j}\,z_{j}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{N_{\mathsf{u}},j}\,z_{j}}=1.$ (A.2)
From independence of $\mathrm{u}_{1},\dots,\mathrm{u}_{N_{\mathsf{u}}}$ assured by Assumption˜6, (A.2) is equivalent to
$\sum_{j=1}^{N_{\mathsf{p}}}m_{i,j}\,z_{j}=0,\quad\forall i=1,\dots,N_{\mathsf{u}},\quad\mbox{i.e.,}\quad\textbf{M}\,\vec{z}=\vec{0},$ (A.3)
for M defined in (20) and $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}$. In other words, any quantity of the form (19) is dimensionless if and only if the vector $\vec{z}$ of exponents in (19) belongs to the null space (kernel) of matrix M (20), i.e., $\vec{z}\in\mathrm{ker}(\textbf{M})$.
Appendix B Proof of Theorem 2
See 2
@@ Appendix B Proof of Theorem 2
First, we show that Assumptions˜4, 5 and 6 imply that the parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21) are dimensionless, strictly positive and independent. Then, we conclude the proof of Theorem 2 by demonstrating that any dimensionless quantity of the form (19) can be written in terms of such $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, as this means that there are no further independent quantities one can consider.
Given the matrix M (20) of exponents in (17) (Assumption˜6), any basis of $\mathrm{ker}(\textbf{M})$ is composed by $N_{\mathsf{0}}\coloneqq\mathrm{dim}(\mathrm{ker}(\textbf{M}))=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ independent vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}\in\mathrm{ker}(\textbf{M})\subseteq\mathbb{R}^{N_{\mathsf{p}}}$. The corresponding $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21) are dimensionless since $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}\in\mathrm{ker}(\textbf{M})$, as guaranteed by Theorem 1 under Assumption˜6.
Then, by employing the notation (1) in (21), one has
@@
$\pi_{k}]=[\tilde{p}_{1}]^{b_{k,1}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{b_{k,{N_{\mathsf{p}}}}}=1,\quad\forall k=1,\dots,N_{\mathsf{0}},$ (B.2)
and $p_{1},\dots,p_{N_{\mathsf{p}}}\in(0,\infty)$ (Assumption˜5), it follows the strict positiveness of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21), i.e.,
@@
Moreover, such $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21) are independent since
@@
On the one hand, $a_{1}=a_{2}=\cdots=a_{N_{\mathsf{0}}}=0$ trivially implies that $\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}}=1$. On the other hand, given $\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}}=1$, it follows from (21) that
@@
which implies, from independence of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ assured by Assumption˜4,
@@
As $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ constitute a basis of $\mathrm{ker}(\textbf{M})$, they are linearly independent vectors. Thus, (B.6) gives $a_{1}=a_{2}=\cdots=a_{N_{\mathsf{0}}}=0$ and (B.4) holds, showing that $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21) are independent.
Finally, any dimensionless quantity of the form $\tilde{p}_{1}^{z_{1}}\,\tilde{p}_{2}^{z_{2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{z_{N_{\mathsf{p}}}}$, with $z_{1},z_{2},\dots,z_{N_{\mathsf{p}}}\in\mathbb{R}$, is such that $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathrm{ker}(\textbf{M})$, as guaranteed by Theorem 1 under Assumption˜6. Then, the vector $\vec{z}$ can be written as a linear combination of the vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ composing a basis of $\mathrm{ker}(\textbf{M})$. By using (21), it follows that
@@
for some $a_{1},\dots,a_{N_{\mathsf{0}}}\in\mathbb{R}$. In conclusion, any dimensionless quantity of the form (19) can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21) only. ∎
Appendix C Proof of Theorem 3
See 3
@@ Appendix C Proof of Theorem 3
As it does for $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21) (see the proof of Theorem 2 in Appendix B), Assumption˜5 ensures that the numerical values $\kappa_{i}$ and $\theta_{j}$ of all $\tilde{\kappa}_{i}$ in (15) and $\tilde{\theta}_{j}$ in (16) are strictly positive. Then, it is possible to plug (15) and (16) into (14) (see Assumptions˜1, 2 and 3). This uncovers that each of the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only, i.e., in the form (19):
$\exists\,\vec{l}_{i}\coloneqq(l_{i,1},l_{i,2},\dots,l_{i,N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}\quad\mbox{such that}\quad\lambda_{i}=\tilde{p}_{1}^{l_{i,1}}\,\tilde{p}_{2}^{l_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{l_{i,N_{\mathsf{p}}}},$ (C.1)
@@
Each of the coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (C.1) is dimensionless and, thus, Theorem 1 guarantees, under Assumption˜6, that each of the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ of exponents in (C.1) belongs to the kernel of the matrix M (20), i.e., $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}\in\mathrm{ker}(\textbf{M})$. Recalling Theorem 2 under Assumptions˜4, 5 and 6, one considers the basis’ vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (21) for $\mathrm{ker}(\textbf{M})$. Then, any vector $\vec{l}_{i}$ (C.1) belonging to $\mathrm{ker}(\textbf{M})$ can be written (in a unique way) as a linear combination of such $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (21), i.e.,
@@
By plugging (C.2) into (C.1) and, then, using (21) (Theorem 2) under Assumptions˜4, 5 and 6, it follows that each of the coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21) only, i.e.,
$\lambda_{i}=\tilde{p}_{1}^{l_{i,1}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{l_{i,N_{\mathsf{p}}}}=\tilde{p}_{1}^{\sum_{k=1}^{N_{\mathsf{0}}}\alpha_{i,k}\,b_{k,1}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\sum_{k=1}^{N_{\mathsf{0}}}\alpha_{i,k}\,b_{k,N_{\mathsf{p}}}}=\pi_{1}^{\alpha_{i,1}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},$ (C.3)
being $l_{i,j}$ and $b_{k,j}$ the $j$-th components of the vectors $\vec{l}_{i}$ and $\vec{b}_{k}$, respectively, for any $i=1,\dots,N_{\mathsf{l}}$, $j=1,\dots,N_{\mathsf{p}}$ and $k=1,\dots,N_{\mathsf{0}}$. We remark that the values of the exponents $\alpha_{i,k}$ in (C.3) are not explicitly known and, in this sense, we state that they are unknowns of scaling procedures. ∎
Appendix D Proof of Theorem 4
See 4